A new car costs ₱450.000.00. This value subsides by 10% each year. 


1
Expert's answer
2021-11-29T01:50:23-0500

1. Write an exponential model that represents this situation after t years.


"f(t)=P(1-r)^t"

"f(t)=450000(1-0.1)^t"

"f(t)=450000(0.9)^t"

2. How much will the car be worth after 5 years?

"f(5)=450000(0.9)^5=265720.50"

The car will be worth ₱265.720.50 after 5 years.


3.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 t & Cost,f(t) \\ \hline
 1 & 405000.00 \\
 \hdashline
 2 & 364500.00 \\
 \hdashline
 3 & 328050.00 \\
 \hdashline
 4 & 295245.00 \\
\end{array}"